OpenStudy (anonymous):

The stage is a right triangle with sides of a=10 and b=20 feet. How long is the other side if it is side c in this formula? (Use a2 + b2 = c2) Round your solution to the nearest tenth. A. 25.6 ft B. 30.0 ft C. 22.4 ft

OpenStudy (anonymous):

C

OpenStudy (anonymous):

10*10= 100 20*20=400 400+100=500 the square root of 500 is 22.4 so C